# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 5. Recognition and uniqueness results for groups and modules
To prove the result it suffices to show that $N_{N_{C_{\mathrm{Aut}(V)}(W)}(B)}(K)$ contains no $p$-elements. Write $K_{1}=C_{K}(V/[V,T])$ and $K_{2}=C_{K}([V,T]/W)$. Then $|K_{1}|=|K_{2}|=q-1$. Let $t\in N_{N_{C_{\mathrm{Aut}(V)}(W)}(B)}(K)$ be a $p$-element so that $[t,K]=1$. Indeed, $t\in N_{C_{\mathrm{Aut}(V)}(W)}(B)$ and $[V,t]$ is normalised by $B$. Since $t$ is a $p$-element, we conclude that $[V,t]\leq[V,T]$. By a similar reasoning, we have that $[V,T,t]\leq W$. Observe that $[[V,T],K_{2},t]=1$ since $t\in C_{\mathrm{Aut}(V)}(W)$ and that $[t,K_{2},[V,T]]=1$. The three subgroups lemma then implies that $[[V,T],t,K_{2}]=1$. Since $B$ commutes with $t$, we have that $[V,T,t]$ is normalised by $T$ and contained in $W$. If $[V,T,t]$ is non-trivial then $C_{W}(T)\cap[V,T,t]\neq 1$ and since $B$ normalises $[V,T,t]$, we deduce that $C_{W}(T)\leq[V,T,t]$ so that $K_{2}$ centralises $C_{W}(T)$. That is, $K_{2}$ centralises both $[V,T]/W$ and $C_{W}(T)$. But $K_{2}\leq N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ and computing in the manner following the discussion after Lemma 2.1, we observe that no element of order $q-1$ acts like this. Hence, $t$ centralises $[V,T]$.
@@
Now, $[V,K_{1},t]=1=[K_{1},t,V]$ and we conclude that $[V,t,K_{1}]=1$. Since $K_{1}$ acts irreducibly on $[V,T]/W$ we deduce that $[V,t]\leq W$ and a similar argument as before implies that $K_{1}$ centralises $C_{W}(T)$ if $[V,t]$ is non-trivial. If $[V,t]$ is non-trivial, then as in the above paragraph, we reduce to checking that no abelian subgroup of $N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ of order $q-1$ centralises both $C_{W}(T)=C_{[V,T,T]}(T)$ and $V/[V,T]$. Once again, following the discussion after Lemma 2.1, there is no such subgroup. Hence, $[V,t]=1$ so that $t=1$. Therefore, we have shown that $N_{\mathrm{Aut}(V)}(KT)=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(B)$ normalises $X$ and $L=X$, as desired. ∎
@@ 6. The potential $\mathcal{F}$-essential subgroups
By the Alperin–Goldschmidt theorem, any saturated fusion system $\mathcal{F}$ on $S$ is determined by $\mathrm{Aut}_{\mathcal{F}}(S)$, together with the collection of all $\mathrm{Aut}_{\mathcal{F}}(P)$ for $P\in\mathcal{E}(\mathcal{F})$. Here we determine the possibilities for such $P\leq S$ when $S\in\{S_{n}(q),S_{\Lambda}(q)\}$. Throughout this section let $q=p^{m}$ be such that $m>1$. Let $n\geq 2$, assume $S=S_{n}(q)$ or $S=S_{\Lambda}(q)$ and write $V$ for the unique elementary abelian subgroup of $S$ of index $q$ (so that either $V=V_{n}(q)$ or $V=\Lambda(q)$ respectively). Let $S_{0}$ be an elementary abelian group of order $q$ such that $S=S_{0}V$, so that $S_{0}$ may be identified with the subgroup $U$ described in Section 2. Set $R:=Z(S)S_{0}$ and $Q:=Z_{2}(S)S_{0}$. We prove the following three results:
@@
Let $\mathcal{F}$ be a saturated fusion system on $S:=S_{n}(q)$ with $12\leqslant n\leqslant p-$ 1. Then either ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ or ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup Q^{S}$. If, in addition, $O_{p}(\mathcal{F})=1$ then one of the following holds:
@@
Let $\mathcal{F}$ be a saturated fusion system on $S:=S_{\Lambda}(q)$. Then ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$. If, in addition, $O_{p}(\mathcal{F})=1$ then ()=$\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$.
@@
Let $S\in\{S_{\Lambda}(q),S_{n}(q)\}$ and $\mathcal{F}$ be a saturated fusion system on $S$ where $12\leqslant n\leqslant p-$ 1 and let $E\in\mathcal{E}(\mathcal{F})$. The following hold:
@@
Suppose that $\mathcal{F}$ is a saturated fusion system on $S$ and $J\leq S$. Assume that $J_{s}<J_{s-1}<\dots<J_{0}=J$ are $\mathrm{Aut}_{\mathcal{F}}(J)$-invariant with $J_{s}\leq\Phi(J)$. If $A\leq\mathrm{Aut}_{S}(J)$ and $[J_{i},A]\leq J_{i+1}$ for $10\leqslant i\leqslant s-$ 1, then $A\leq O_{p}(\mathrm{Aut}_{\mathcal{F}}(J))$. In particular, if $A\not\leq\mathrm{Inn}(J)$, then $J$ is not $\mathcal{F}$-centric-radical.
@@
Let $\mathcal{F}$ be a saturated fusion system on $S\cong S_{p}(q)$. Then ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}$ and $V\trianglelefteq\mathcal{F}$.
@@
a contradiction by Lemma 6.4. Hence, $V=E$ and ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}$. Since $V$ is characteristic in $S$, $V\trianglelefteq\mathcal{F}$ by [AKO11, Proposition I.4.5]. ∎
@@
We now embark on an initial determination of the essential subgroups of $\mathcal{F}$, where $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ and $12\leqslant n\leqslant p-$ 1. Recall from Lemmas 2.4 and 2.6 that for $S=S_{n}(q)$ we have
$[S\colon V]=q,\hskip 5.69054pt|Z_{i}(S)|=q^{i}\mbox{ for all }1\leqslant i\leqslant n\hskip 5.69054pt\mbox{ and }\hskip 5.69054ptZ_{2}(S)\leq Z_{n}(S)=S^{\prime}=[V,S]<V.$
@@
$[S\colon V]=q,\hskip 5.69054pt|Z_{i}(S)|=q^{i+1}\mbox{ for all }1\leqslant i\leqslant p-1,\hskip 5.69054pt\mbox{ and }\hskip 5.69054ptZ_{2}(S)\leq Z_{p-1}(S)=S^{\prime}=[V,S]<V.$
The proof of the forthcoming lemma follows the same methodology as Proposition 6.5 but reveals the subgroups $R$ and $Q$ as potential candidates for the $\mathcal{F}$-essential subgroups ( in (1) and (2) of Lemma 6.7 respectively). Recall the notation $\mathcal{B}(S)$ and $\mathcal{C}(S)$ introduced before Lemma 2.17 and that $d=q$ if $S=S_{\Lambda}(q)$ and $d=1$ if $S=S_{n}(q)$.
@@
Let $\mathcal{F}$ be a fusion system on $S$ where $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ and $12\leqslant n\leqslant p-$ 1. If $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$, then either
@@
Since $E$ is $\mathcal{F}$-centric, we have $E\nleq V$ and $[Z_{2}(S),E]\leq[Z_{2}(S),S]\leq Z(S)\leq E$, so $Z_{2}(S)\leq N_{S}(E)$. We will freely use the results from Lemma 2.4 and Lemma 2.12 throughout this proof.
@@
and so $Z(E)=Z(S)$. Since $[Z_{2}(S),E]\leq Z(E)$, $Z_{2}(S)$ centralises the sequence $1\trianglelefteq Z(E)\trianglelefteq E$, and we conclude from Lemma 6.4 that $Z_{2}(S)\leq E$. In particular, $[Z_{3}(S),E]\leq Z_{2}(S)\leq E$ and $Z_{3}(S)\leq N_{S}(E)$. Since $E\cap V$ is an elementary abelian subgroup of $E$ containing $Z_{2}(S)$, $|E\cap V|\geqslant q^{2}d$. Moreover $N_{V}(E)$ centralises the normal chain $1\trianglelefteq E\cap V\trianglelefteq E$ and so Lemma 6.4 implies that either $V\leq E$, or $E\cap V$ is not $\mathrm{Aut}_{\mathcal{F}}(E)$-invariant.
If $V\leq E$ then $V$ is the unique elementary abelian subgroup of $E$ of maximal order. Moreover, if $A$ is such that $V<A\leq E$, then $Z_{i}(A)=Z_{i}(S)$ for $1\leqslant i\leqslant m:=\mathrm{min}\{n,p-1\}$ and $[S,A]\leq Z_{m}(S)$. Then $S$ centralises the series $1\trianglelefteq Z(E)\trianglelefteq Z_{2}(E)\trianglelefteq\dots\trianglelefteq Z_{m}(E)\trianglelefteq E$ and Lemma 6.4 yields a contradiction.
If $E\cap V$ is not $\mathrm{Aut}_{\mathcal{F}}(E)$-invariant then there exists an elementary abelian subgroup $A$ of $E$ distinct from $E\cap V$ with $|A|\geqslant|E\cap V|\geqslant q^{2}d$. On the other hand if $a\in A\setminus V$ then $A\cap V\leq C_{V}(a)=C_{V}(S)\leq Z(S)$ which implies that $|A\cap V|\leqslant qd$ so $|A|=|AV/V||A\cap V|\leqslant q^{2}d$. Therefore $|E\cap V|=|A|=q^{2}d$ and we we must have $S=AV=EV$ and $Z_{2}(S)=E\cap V$. It follows that $E=AZ_{2}(S)$ has order $q^{3}d$ and $\Phi(E)=[A,Z_{2}(S)]\leq Z(S)=Z(E)$. Moreover $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))$ centralises $Z(E)$ and $N_{S}(E)$ centralises $Z_{2}(S)/Z(E)$. Since $|E/Z(E):Z_{2}(S)/Z(E)|=q$, we obtain $q=|Z_{3}(S)E/E|\leqslant|N_{S}(E)/E|$ and Lemma 5.5 implies that $N_{S}(E)=Z_{3}(S)E$ and $E/Z(E)$ may be regarded as a natural $\operatorname{SL}_{2}(q)$-module for $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(E))\cong\operatorname{SL}_{2}(q)$. This completes the proof of (2). ∎
@@
Suppose that $E\in\mathcal{E}(\mathcal{F})$. Then $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))}(\mathrm{Aut}_{S}(E))$ has cyclic Hall $p^{\prime}$-subgroups of order $q-1$. Fix such a subgroup $T_{E}$, and let $T\leq\mathrm{Aut}_{\mathcal{F}}(S)$ be a subgroup of order $q-1$ such that $T|_{E}=T_{E}$. Then
@@
Let $t\in T^{\#}$. By Lemma 6.7, $E\cap V\in\{Z(S),Z_{2}(S)\}$ and so $E\cap V$ is normalised by $t$. Since $S=EV$, the elements of $(S/V)^{\#}$ are of the form $eV$ where $e\in E\setminus(E\cap V)$. Suppose $[t,eV]=1$ for some such $e$. Then $[t,e]\leq E\cap V$, and since $E/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E)))$ is a natural $\operatorname{SL}_{2}(q)$-module with
$[E/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))),N_{S}(E)]=(E\cap V)/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))),$
@@
Let $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$. We first show that $[V,S]E=[V,S]S_{0}$ and $E$ is $S$-conjugate to either $R$ or $Q$.
Let $T_{E}$ and $T$ be as in Lemma 6.11 and recall the definitions of $\Sigma$ and $\delta$ from Section 2. Choose $\tau_{E}\in T$ an element of order $q-1$ and observe that $\tau_{E}$ is conjugate in $\mathrm{Aut}(S)$ to an element of $\mathrm{Aut}_{P_{n}^{*}(q)}(S)$. Since the isomorphism type of $\mathcal{F}$ is unchanged by conjugating by elements of $\mathrm{Aut}(S)$, we may assume that $\tau_{E}\in\mathrm{Aut}_{P_{n}^{*}(q)}(S)$ and ${\tau_{E}}|_{E}$ is a generator of $T_{E}$.
Suppose first that $E$ is abelian. Recalling the definition of $\delta$ which precedes Lemma 2.20 we see that $\tau_{E}\delta=(\tau,\tau^{-1})$ where $\tau$ is some primitive element of $\mathbb{K}$. We know that $\tau_{E}$ normalises $V$, $[V,S]S_{0}$ and $[V,S]E$.
Recall the definition of the group $\Sigma$ from Section 2.
@@
Now, $\tau_{E}$ acts on $S/[V,S]$ as $\mathrm{diag}(\tau^{-1},\tau^{n+1})$ preserving the subgroups $V/[V,S]$ and $S_{0}[V,S]/[V,S]$. Assume that $E[V,S]\neq S_{0}[V,S]$. Then, as $\mathbb{F}_{p}\langle\tau_{E}\rangle$-modules, we must have that $V/[V,S]$ and $[V,S]S_{0}/[V,S]$ are isomorphic. This means that
@@
()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}\cup Q^{S}.$
Assume that some $S$-conjugate of $R$ is $\mathcal{F}$-essential. Then all $S$-conjugates of $R$ are $\mathcal{F}$-essential. In particular, $R\in\mathcal{E}(\mathcal{F})$. Similarly, if some $S$-conjugate of $Q$ is $\mathcal{F}$-essential then $Q\in\mathcal{E}(\mathcal{F})$. Note that $R<Q$. Hence by Lemma 6.8 we ascertain that either ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ or else ()$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup Q^{S}$.
Finally, suppose that $O_{p}(\mathcal{F})=1$. Applying [AKO11, Proposition I.4.5] we see that ()$\mathcal{E}(\mathcal{F})\neq\{V\}$ since $V$ is characteristic in $S$. If ()=$\mathcal{E}(\mathcal{F})=\{Q^{S}\}$ then $Z(S)$ is left invariant by $\mathrm{Aut}_{\mathcal{F}}(S)$ and $\mathrm{Aut}_{\mathcal{F}}(F)$ for any $F\in Q^{S}$ and so $Z(S)\trianglelefteq\mathcal{F}$ by [AKO11, Proposition I.4.5], a contradiction. Thus ()$\mathcal{E}(\mathcal{F})\neq\{Q^{S}\}$, as required. ∎

### v2

@@ 5. Recognition and uniqueness results for groups and modules
To prove the result it suffices to show that $N_{N_{C_{\mathrm{Aut}(V)}(W)}(TK)}(K)$ contains no $p$-elements. Write $K_{1}=C_{K}(V/[V,T])$ and $K_{2}=C_{K}([V,T]/W)$. Then $|K_{1}|=|K_{2}|=q-1$. Let $t\in N_{N_{C_{\mathrm{Aut}(V)}(W)}(TK)}(K)$ be a $p$-element so that $[t,K]=1$. Indeed, $t\in N_{C_{\mathrm{Aut}(V)}(W)}(TK)$ and $[V,t]$ is normalised by $TK$. Since $t$ is a $p$-element, we conclude that $V,t]\leq[V,T]$. By a similar reasoning, we have that $V,T,t]\leq W$. Observe that $V,T],K_{2},t]=1$ since $t\in C_{\mathrm{Aut}(V)}(W)$ and that $t,K_{2},[V,T]]=1$. The three subgroups lemma then implies that $V,T],t,K_{2}]=1$. Since $TK$ commutes with $t$, we have that $[V,T,t]$ is normalised by $T$ and contained in $W$. If $[V,T,t]$ is non-trivial then $C_{W}(T)\cap[V,T,t]\neq 1$ and since $TK$ normalises $[V,T,t]$, we deduce that $C_{W}(T)\leq[V,T,t]$ so that $K_{2}$ centralises $C_{W}(T)$. That is, $K_{2}$ centralises both $[V,T]/W$ and $C_{W}(T)$. But $K_{2}\leq N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ and computing in the manner following the discussion after Lemma 2.4, we observe that no element of order q1q-1 acts like this. Hence, $t$ centralises $[V,T]$.
@@
Now, $V,K_{1},t]=1=[K_{1},t,V]$ and we conclude that $V,t,K_{1}]=1$. Since $K_{1}$ acts irreducibly on $[V,T]/W$ we deduce that $V,t]\leq W$ and a similar argument as before implies that $K_{1}$ centralises $C_{W}(T)$ if $[V,t]$ is non-trivial. If $[V,t]$ is non-trivial, then as in the above paragraph, we reduce to checking that no abelian subgroup of $N_{\mathrm{Aut}_{D^{*}}(V)}(T)$ of order q1q-1 centralises both $C_{W}(T)=C_{[V,T,T]}(T)$ and $V/[V,T]$. Once again, following the discussion after Lemma 2.4, there is no such subgroup. Hence, $[V,t]=1$ so that $t=1$. Therefore, we have shown that $N_{\mathrm{Aut}(V)}(KT)=N_{\mathrm{Aut}_{D^{\dagger}}(V)}(TK)$ normalises $X$ and $L=X$, as desired. ∎
@@ 6. The potential $\mathcal{F}$-essential subgroups
By the Alperin–Goldschmidt theorem, any saturated fusion system $\mathcal{F}$ on $S$ is determined by $\mathrm{Aut}_{\mathcal{F}}(S)$, together with the collection of all $\mathrm{Aut}_{\mathcal{F}}(P)$ for $P\in\mathcal{E}(\mathcal{F})$. Here we determine the possibilities for such $P\leq S$ when $S\in\{S_{n}(q),S_{\Lambda}(q)\}$. Throughout this section let $q=p^{m}$ be such that $m>1$. Let $n\geq 2$, assume $S=S_{n}(q)$ or $S=S_{\Lambda}(q)$ and write $V$ for the unique elementary abelian subgroup of $S$ of index $q$ (so that either $V=V_{n}(q)$ or $V=\Lambda(q)$ respectively). Let $S_{0}$ be an elementary abelian group of order $q$ such that $S=S_{0}V$, so that $S_{0}$ may be identified with the subgroup $U$ described from Notation 2.2. Recall $R:=Z(S)S_{0}$ and $Q:=Z_{2}(S)S_{0}$ from Notation 2.21. We prove the following three results:
@@
Let $\mathcal{F}$ be a saturated fusion system on $S:=S_{n}(q)$ with $2\leqslant n\leqslant p-1$. Then either $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ or $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup Q^{S}$. If, in addition, $O_{p}(\mathcal{F})=1$ then one of the following holds:
@@
Let $\mathcal{F}$ be a saturated fusion system on $S:=S_{\Lambda}(q)$. Then $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$. If, in addition, $O_{p}(\mathcal{F})=1$ then $\mathcal{E}(\mathcal{F})=\{V\}\cup R^{S}$.
@@
Let $S\in\{S_{\Lambda}(q),S_{n}(q)\}$ and $\mathcal{F}$ be a saturated fusion system on $S$ where $2\leqslant n\leqslant p-1$ and let $E\in\mathcal{E}(\mathcal{F})$. The following hold:
@@
Suppose that $\mathcal{F}$ is a saturated fusion system on $S$ and $J\leq S$. Assume that $J_{s}<J_{s-1}<\dots<J_{0}=J$ are $\mathrm{Aut}_{\mathcal{F}}(J)$-invariant with $J_{s}\leq\Phi(J)$. If $A\leq\mathrm{Aut}_{S}(J)$ and $J_{i},A]\leq J_{i+1}$ for $0\leqslant i\leqslant s-1$, then $A\leq O_{p}(\mathrm{Aut}_{\mathcal{F}}(J))$. In particular, if $A\not\leq\mathrm{Inn}(J)$, then $J$ is not $\mathcal{F}$-centric-radical.
@@
Let $\mathcal{F}$ be a saturated fusion system on $S\cong S_{p}(q)$. Then $\mathcal{E}(\mathcal{F})\subseteq\{V\}$ and $V\trianglelefteq\mathcal{F}$.
@@
a contradiction by Lemma 6.4. Hence, $V=E$ and $\mathcal{E}(\mathcal{F})\subseteq\{V\}$. Since $V$ is characteristic in $S$, $V\trianglelefteq\mathcal{F}$ by [AKO11, Proposition I.4.5]. ∎
@@
We now embark on an initial determination of the essential subgroups of $\mathcal{F}$, where $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ and $2\leqslant n\leqslant p-1$. Recall from Lemmas 2.8 and 2.10 that for $S=S_{n}(q)$ we have
$S\colon V]=q,\hskip 5.69054pt|Z_{i}(S)|=q^{i}\mbox{ for all }1\leqslant i\leqslant n\hskip 5.69054pt\mbox{ and }\hskip 5.69054ptZ_{2}(S)\leq Z_{n}(S)=S^{\prime}=[V,S]<V.$
@@
$S\colon V]=q,\hskip 5.69054pt|Z_{i}(S)|=q^{i+1}\mbox{ for all }1\leqslant i\leqslant p-1,\hskip 5.69054pt\mbox{ and }\hskip 5.69054ptZ_{2}(S)\leq Z_{p-1}(S)=S^{\prime}=[V,S]<V.$
The proof of the forthcoming lemma follows the same methodology as Proposition 6.5 but reveals the subgroups $R$ and $Q$ as potential candidates for the $\mathcal{F}$-essential subgroups (in (1) and (2) of Lemma 6.7 respectively). Recall the notation $\mathcal{B}(S)$ and $\mathcal{C}(S)$ introduced in Notation 2.21 and that $d=q$ if $S=S_{\Lambda}(q)$ and $d=1$ if $S=S_{n}(q)$.
@@
Let $\mathcal{F}$ be a fusion system on $S$ where $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ and $2\leqslant n\leqslant p-1$. If $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$, then either
@@
Since $E$ is $\mathcal{F}$-centric, we have $E\nleq V$ and $Z_{2}(S),E]\leq[Z_{2}(S),S]\leq Z(S)\leq E$, so $Z_{2}(S)\leq N_{S}(E)$. We will freely use the results from Lemma 2.8 and Lemma 2.16 throughout this proof.
@@
and so $Z(E)=Z(S)$. Since $Z_{2}(S),E]\leq Z(E)$, $Z_{2}(S)$ centralises the sequence $1\trianglelefteq Z(E)\trianglelefteq E$, and we conclude from Lemma 6.4 that $Z_{2}(S)\leq E$. In particular, $Z_{3}(S),E]\leq Z_{2}(S)\leq E$ and $Z_{3}(S)\leq N_{S}(E)$. Since $E\cap V$ is an elementary abelian subgroup of $E$ containing $Z_{2}(S)$, $|E\cap V|\geqslant q^{2}d$. Moreover $N_{V}(E)$ centralises the normal chain $1\trianglelefteq E\cap V\trianglelefteq E$ and so Lemma 6.4 implies that either $V\leq E$, or $E\cap V$ is not $\mathrm{Aut}_{\mathcal{F}}(E)$-invariant.
If $V\leq E$ then $V$ is the unique elementary abelian subgroup of $E$ of maximal order. Moreover, if $A$ is such that $V<A\leq E$, then $Z_{i}(A)=Z_{i}(S)$ for $1\leqslant i\leqslant m:=\mathrm{min}\{n,p-1\}$ and $S,A]\leq Z_{m}(S)$. Then $S$ centralises the series $1\trianglelefteq Z(E)\trianglelefteq Z_{2}(E)\trianglelefteq\dots\trianglelefteq Z_{m}(E)\trianglelefteq E$ and Lemma 6.4 yields a contradiction.
If $E\cap V$ is not $\mathrm{Aut}_{\mathcal{F}}(E)$-invariant then there exists an elementary abelian subgroup $A$ of $E$ distinct from $E\cap V$ with $|A|\geqslant|E\cap V|\geqslant q^{2}d$. On the other hand if $a\in A\setminus V$ then $A\cap V\leq C_{V}(a)=C_{V}(S)\leq Z(S)$ which implies that $|A\cap V|\leqslant qd$ so $|A|=|AV/V||A\cap V|\leqslant q^{2}d$. Therefore $|E\cap V|=|A|=q^{2}d$ and we must have $S=AV=EV$ and $Z_{2}(S)=E\cap V$. It follows that $E=AZ_{2}(S)$ has order $q^{3}d$ and $\Phi(E)=[A,Z_{2}(S)]\leq Z(S)=Z(E)$. Moreover $O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))$ centralises $Z(E)$ and $N_{S}(E)$ centralises $Z_{2}(S)/Z(E)$. Since $|E/Z(E):Z_{2}(S)/Z(E)|=q$, we obtain $q=|Z_{3}(S)E/E|\leqslant|N_{S}(E)/E|$ and Lemma 5.5 implies that $N_{S}(E)=Z_{3}(S)E$ and $E/Z(E)$ may be regarded as a natural $\operatorname{SL}_{2}(q)$-module for $O^{p^{\prime}}(\mathrm{Out}_{\mathcal{F}}(E))\cong\operatorname{SL}_{2}(q)$. This completes the proof of (2). ∎
@@
Suppose that $E\in\mathcal{E}(\mathcal{F})$. Then $N_{O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))}(\mathrm{Aut}_{S}(E))$ has cyclic Hall $p^{\prime}$-subgroups of order q1q-1. Fix such a subgroup $T_{E}$, and let $T\leq\mathrm{Aut}_{\mathcal{F}}(S)$ be a subgroup of order q1q-1 such that $T|_{E}=T_{E}$. Then
@@
Let $t\in T^{\#}$. By Lemma 6.7, $E\cap V\in\{Z(S),Z_{2}(S)\}$ and so $E\cap V$ is normalised by $t$. Since $S=EV$, the elements of $(S/V)^{\#}$ are of the form $eV$ where $e\in E\setminus(E\cap V)$. Suppose $[t,eV]=1$ for some such $e$. Then $t,e]\leq E\cap V$, and since $E/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E)))$ is a natural $\operatorname{SL}_{2}(q)$-module with
$E/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))),N_{S}(E)]=(E\cap V)/C_{E}(O^{p^{\prime}}(\mathrm{Aut}_{\mathcal{F}}(E))),$
@@
Let $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$. We first show that $V,S]E=[V,S]S_{0}$ and $E$ is $S$-conjugate to either $R$ or $Q$.
Let $T_{E}$ and $T$ be as in Lemma 6.11 and recall the definitions of $\Sigma$ from Notation 2.2, $P_{n}^{*}(q)$ from Notation 2.3 and $\delta$ from Notation 2.25. Choose $\tau_{E}\in T$ an element of order q1q-1 and observe that $\tau_{E}$ is conjugate in $\mathrm{Aut}(S)$ to an element of $\mathrm{Aut}_{P_{n}^{*}(q)}(S)$. Since the isomorphism type of $\mathcal{F}$ is unchanged by conjugating by elements of $\mathrm{Aut}(S)$, we may assume that $\tau_{E}\in\mathrm{Aut}_{P_{n}^{*}(q)}(S)$ and ${\tau_{E}}|_{E}$ is a generator of $T_{E}$.
Suppose first that $E$ is abelian. We see that $\tau_{E}\delta=(\tau,\tau^{-1})$ where $\tau$ is some primitive element of $\mathbb{K}$. We know that $\tau_{E}$ normalises $V$, $V,S]S_{0}$ and $[V,S]E$.
@@
Now, $\tau_{E}$ acts on $S/[V,S]$ as $\mathrm{diag}(\tau^{-1},\tau^{n+1})$ preserving the subgroups $V/[V,S]$ and $S_{0}[V,S]/[V,S]$. Assume that $E[V,S]\neq S_{0}[V,S]$. Then, as $\mathbb{F}_{p}\langle\tau_{E}\rangle$-modules, we must have that $V/[V,S]$ and $V,S]S_{0}/[V,S]$ are isomorphic. This means that
@@
$\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}\cup Q^{S}.$
Assume that some $S$-conjugate of $R$ is $\mathcal{F}$-essential. Then all $S$-conjugates of $R$ are $\mathcal{F}$-essential. In particular, $R\in\mathcal{E}(\mathcal{F})$. Similarly, if some $S$-conjugate of $Q$ is $\mathcal{F}$-essential then $Q\in\mathcal{E}(\mathcal{F})$. Note that $R<Q$. Hence by Lemma 6.8 we ascertain that either $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup R^{S}$ or else $\mathcal{E}(\mathcal{F})\subseteq\{V\}\cup Q^{S}$.
Finally, suppose that $O_{p}(\mathcal{F})=1$. Applying [AKO11, Proposition I.4.5] we see that $\mathcal{E}(\mathcal{F})\neq\{V\}$ since $V$ is characteristic in $S$. If $\mathcal{E}(\mathcal{F})=\{Q^{S}\}$ then $Z(S)$ is left invariant by $\mathrm{Aut}_{\mathcal{F}}(S)$ and $\mathrm{Aut}_{\mathcal{F}}(F)$ for any $F\in Q^{S}$ and so $Z(S)\trianglelefteq\mathcal{F}$ by [AKO11, Proposition I.4.5], a contradiction. Thus $\mathcal{E}(\mathcal{F})\neq\{Q^{S}\}$, as required. ∎